Show that the thermal efficiency of the Carnot cycle in terms of the isentropic compression ratio π‘Ÿπ‘˜ is given by 𝑒 = 1 βˆ’ (1 /(π‘Ÿπ‘˜) π‘˜βˆ’1).Β 


1
Expert's answer
2021-11-25T11:51:18-0500

thermal efficiency:

"𝑒 = 1 − \frac{1 }{𝑟_𝑘^{𝑘−1}}"


"e=\frac{work\ done}{heat\ supplied}"


for Isentropic Compression 4-1:

"e=1-\frac{T_3}{T_2}"


"\frac{T_3}{T_2}=(\frac{V_2}{V_3})^{k-1}=(\frac{1}{r_k})^{k-1}"


"𝑒 = 1 − \frac{1 }{𝑟_𝑘^{𝑘−1}}"